OpenStudy (amina_soneviseth):

MEDALS GIVEN what is the reciprecal of 7/2a? (its a fraction)

OpenStudy (jadedry):

The reciprocal of a fraction is its "inverse" . When the reciprocal is multiplied by the original number, the product will always be one. Therefore the reciprocal of \[\frac{ 7 }{ 2a} = \frac{ 2a }{7}\]

OpenStudy (jhannybean):

apply this rule to your fraction: \(\dfrac{1}{\dfrac{a}{b}} = \dfrac{b}{a}\) where \(\dfrac{a}{b}\) is your fraction.

OpenStudy (amina_soneviseth):

@Jadedry thank you very much!

OpenStudy (jadedry):

You're welcome. c: